# Agent-based modeling for realistic reproduction of human mobility and contact behavior to evaluate test and isolation strategies in epidemic infectious disease spread

### v1

@@ 3.3 Effects from combined variation of testing and isolation parameters
Figure 9: Aggregated outcomes for a variation of voluntary likelihood for symptomatic agents to test, $p_{\textrm{s}}$, against ratio for nonsymptomatic to symptomatic agents to test, $1/\mu_{\textrm{ns}}$. Aggregated outcomes on cumulative infections (top left), deaths (top right), maximum hospitalized patients (bottom left), and maximum daily infections (bottom right). Testing likelihood for symptomatic patients increases from top to bottom. Ratio for testing of symptomatic against symptomatic increases from left to right.
Substantial changes in different parameters such as the likelihood for symptomatic testing naturally induce changed outcomes of our simulation. However, varying single parameters is often insufficient to get more insight into the model as combined effects are then overlooked. In this section, we want to explore the combined effects of different testing and isolation parameters. We present all combined effects with respect to four different outcomes: the cumulative number of infections, the number of deaths, the maximum number of hospitalized patients and the maximum number of daily infections. While it is clear why to consider the first two indicators, maximum numbers are of high importance as capacities of the health sector should not be attained.
@@
Figure 10: Aggregated outcomes for a variation of voluntary likelihood for symptomatic agents to test, $p_{\textrm{s}}$, against quarantine length, $q_{\textrm{d}}$. Aggregated outcomes on cumulative infections (top left), deaths (top right), maximum hospitalized patients (bottom left), and maximum daily infections (bottom right). Testing likelihood for symptomatic patients increases from top to bottom. Quarantine length in number of days increases from left to right.
@@
Figure 11: Aggregated outcomes for a variation of quarantine length, $q_{\textrm{d}}$, against quarantine efficiency, $q_{\textrm{e}}$. Aggregated outcomes on cumulative infections (top left), deaths (top right), maximum hospitalized patients (bottom left), and maximum daily infections (bottom right). Quarantine length in number of days increases from top to bottom. Quarantine efficiency $q_{e}$ increases from left to right.
@@ 3.4 Increased symptomatic testing
Figure 12: Impact of increased symptomatic testing. In our ABM fitted to data from Brunswick between March 1st and May 30, 2021 (blue), the symptom-based testing probability was doubled (orange) and tripled (green). The lockdown was active from end of March until end of April, while test frequency was further reduced in the Easter week.
@@ 3.5 Counterfactual scenarios with increased testing instead of lockdowns
Figure 13: Counterfactual scenario without lockdown but increased symptom-based testing likelihood in Brunswick between March 1st and May 30, 2021. Standard fitted scenario without lockdown (blue) against a three- (orange) and fivefold (green) symptom-based testing rate. During the Easter week, testing was reduced.
@@ 4 Discussion
Although ABMs allow a very detailed representation of the real world, still simplifications and assumptions have to be made. While we believe to have integrated the most important features for the considered time period, additional features or extensions might be needed for the consideration of novel virus variants or long-term waning immunity. Outcomes of the presented study have been obtained in the light of a mostly immune-naive population and outcomes could be different for a highly immunized population, late-phase epidemic or endemic scenarios. Furthermore, in this study, we have considered neither the (short-term) protection against transmission, nor damped viral load courses, through vaccinations or previous infections. These effects should be integrated in an analysis for later stages of the pandemic.
@@
Our findings suggest that quarantine length seems to play a more important role than quarantine efficiency and a minimum quarantine of 5-8 days appears to be inevitable for quarantine to have a substantial effect. While, in the literature, it is stated that even a 14-day quarantine cannot capture all uncommon or outlier infections [84], several authors obtained similar findings [83, 80, 85, 77, 86] with respect to quarantine length and differently optimized quarantining strategies. In [85], the authors studied different test-and-release strategies. They found an diminishing relative utility of quarantine relative to a 10-day strategy and almost no additional benefit for quarantines beyond 10 days. In [80], the authors found a five-day quarantine quite efficient, with increased effect when testing is more widespread (compare to  Fig. 10). In [83], the authors naturally observed increased effects for quarantines of 7, 14, and 21 days but also found that less than five percent of cases are missed with a 7-day quarantine and different test-exit strategies. The authors of [87] found 5-10 days of quarantine most important for preventing transmissions and highlighted how test-and-release based quarantining with shorter periods can outperform 14-days quarantining with high compliance. In [86], the authors found that testing on exit can reduce a 14-day quarantine by 50 %.
@@ 5 Conclusion
In this study, we presented an efficiently implemented and parallelized ABM that can be executed with hundreds of thousands to millions of agents on a consumer laptop. Furthermore, it allows the use of a hybrid, shared and distributed memory parallelism to compute large numbers of ensemble results on a supercomputing facility. This capability is essential for handling the computational complexity of ABMs, especially when exploring wide parameter space or performing sensitivity analysis.
@@
We furthermore studied the combined effects of symptom-based and untargeted, symptom-independent testing strategies as well as quarantine length and efficiency. We observed that with increased testing of symptomatic persons, the testing of nonsymptomatic persons plays a higher role, leading to a larger reduction in all considered endpoints. On the other hand, with weak symptomatic control, i.e., with a low likelihood to test symptomatic persons, untargeted testing was observed to be mostly inefficient. Furthermore, quarantine length seems to play a more important role than quarantine efficiency and a minimum quarantine of 5-8 days appears to be inevitable for quarantine to have a substantial effect. This, of course, is in line with high viral loads that have been observed at the beginning of the infection. Similar findings have been discussed in the previous section. Finally, we also observed how high symptomatic testing rates leverage quarantining effects such that even short quarantining can have a large effect on mitigation.
@@ Acknowledgments
The authors gratefully acknowledge computing time on the supercomputer JURECA [63] at Forschungszentrum Jülich under grant HPC4EPI. DK, MMH, and MJK have received funding from the Initiative and Networking Fund of the Helmholtz Association (grant agreement number KA1-Co-08, Project LOKI-Pandemics). The authors DA, CG, AS, and MJK have received funding by the German Federal Ministry for Digital and Transport under grant agreement FKZ19F2211A (Project PANDEMOS), KN has received funding by the German Federal Ministry for Digital and Transport under grant agreement FKZ19F2211B (Project PANDEMOS). SK has performed his work as part of the Helmholtz School for Data Science in Life, Earth and Energy (HDS-LEE) and received funding from the Helmholtz Association of German Research Centres. The funders had no role in study design, data collection and analysis, decision to publish, or preparation of the manuscript.

### v2

@@ 3.3 Effects from combined variation of testing and isolation parameters
Figure 9: Aggregated outcomes for a variation of voluntary likelihood for symptomatic agents to test, $p_{\textrm{s}}$, against ratio for nonsymptomatic to symptomatic agents to test, $1/\mu_{\textrm{ns}}$. Aggregated outcomes on cumulative infections (top left), deaths (top right), maximum hospitalized patients (bottom left), and maximum hourly infections (bottom right). Testing likelihood for symptomatic patients increases from top to bottom. Ratio for testing of symptomatic against symptomatic increases from left to right.
Substantial changes in different parameters such as the likelihood for symptomatic testing naturally induce changed outcomes of our simulation. However, varying single parameters is often insufficient to get more insight into the model as combined effects are then overlooked. In this section, we want to explore the combined effects of different testing and isolation parameters. We present all combined effects with respect to four different outcomes: the cumulative number of infections, the number of deaths, the maximum number of hospitalized patients and the maximum number of hourly infections. While it is clear why to consider the first two indicators, maximum numbers are of high importance as capacities of the health sector should not be attained.
@@
Figure 10: Aggregated outcomes for a variation of voluntary likelihood for symptomatic agents to test, $p_{\textrm{s}}$, against quarantine length, $q_{\textrm{d}}$. Aggregated outcomes on cumulative infections (top left), deaths (top right), maximum hospitalized patients (bottom left), and maximum hourly infections (bottom right). Testing likelihood for symptomatic patients increases from top to bottom. Quarantine length in number of days increases from left to right.
@@
Figure 11: Aggregated outcomes for a variation of quarantine length, $q_{\textrm{d}}$, against quarantine efficiency, $q_{\textrm{e}}$. Aggregated outcomes on cumulative infections (top left), deaths (top right), maximum hospitalized patients (bottom left), and maximum hourly infections (bottom right). Quarantine length in number of days increases from top to bottom. Quarantine efficiency $q_{e}$ increases from left to right.
@@ 3.4 Increased symptomatic testing
Figure 12: Impact of increased symptomatic testing. In our ABM fitted to data from Brunswick between March 1st and May 30, 2021 (blue), the symptom-based testing probability was doubled (orange) and tripled (green), where the shaded area represents the 50% confidence interval. The lockdown was active from end of March until end of April, while test frequency was further reduced in the Easter week, as indicated by vertical lines. Furthermore, there are special Easter gatherings on Easter Sunday and Monday.
@@ 3.5 Counterfactual scenarios with increased testing instead of lockdowns
Figure 13: Counterfactual scenario without lockdown but increased symptom-based testing likelihood in Brunswick between March 1st and May 30, 2021. Standard fitted scenario without lockdown (blue) against a three- (orange) and fivefold (green) symptom-based testing rate, where the shaded area represents the 50% confidence interval. During the Easter week, testing was reduced, as indicated by vertical lines. Furthermore, there are special Easter gatherings on Easter Sunday and Monday.
@@ 4 Discussion
As a natural limitation, it is unclear to what extent the presented outcomes can be transferred to regions with settings which differ in mobility patterns, healthcare systems, and cultural practices from the Brunswick region. For Germany, we, e.g., obtained similar mobility patterns and trips for the regions of Brunswick and Munich, although both are of substantially different size and located within a different geographical structure, compare Figure 14 to Figure 18. For regions with substantially different contact patterns or activities, our findings might not hold true. However, our framework provides the possibility to insert generic contact and trip patterns to analyze scenarios in other regions where these patterns deviate. Enhanced testing efforts may not be as critical for controlling infection spread if the daily number of contacts is substantially smaller. A general result cannot be inferred from our findings, as these results highly depend on the contact structure (minimum, average, maximum or distribution) and on the type and intensity of contacts. These properties depend, in particular, on the cultural and socioeconomic contexts. In addition, for areas with limited healthcare infrastructure (e.g., a low number of available hospital beds), high testing rates can serve as an effective, minimally invasive intervention to curb infection spread and prevent overwhelming the healthcare system. Furthermore, with the ABM presented here being itself a result, the openly available source code offers scientists the capability of integrating and exploring arbitrary populations and mobility patterns to infer potentially changing outcomes.
@@
Furthermore, although ABMs allow a very detailed representation of the real world, simplifications and assumptions still have to be made. While we believe to have integrated the most important features for the considered time period, additional features or extensions might be needed for the consideration of novel virus variants or long-term waning immunity. Outcomes of the presented study have been obtained in the light of a mostly immune-naive population and outcomes could be different for a highly immunized population, late-phase epidemic or endemic scenarios. Additionally, in this study, we have considered neither the (short-term) protection against transmission, nor damped viral load courses, through vaccinations or previous infections. These effects should be integrated in an analysis for later stages of the pandemic.
@@
Our findings suggest that quarantine length seems to play a more important role than quarantine efficiency and a minimum quarantine of 5–8 days appears to be inevitable for quarantine to have a substantial effect. While, in the literature, it is stated that even a 14-day quarantine cannot capture all uncommon or outlier infections [94], several authors obtained similar findings [93, 90, 95, 87, 96] with respect to quarantine length and differently optimized quarantining strategies. In [95], the authors studied different test-and-release strategies. They found an diminishing relative utility of quarantine relative to a 10-day strategy and almost no additional benefit for quarantines beyond 10 days. In [90], the authors found a five-day quarantine quite efficient, with increased effect when testing is more widespread (compare to Fig. 10). In [93], the authors naturally observed increased effects for quarantines of 7, 14, and 21 days but also found that less than five percent of cases are missed with a 7-day quarantine and different test-exit strategies. The authors of [97] found 5–10 days of quarantine most important for preventing transmissions and highlighted how test-and-release based quarantining with shorter periods can outperform 14-days quarantining with high compliance. In [96], the authors found that testing on exit can reduce a 14-day quarantine by 50 %.
The combination of, both, frequent tests, in particular, for symptomatic persons, and adequate quarantine shows an effective way to minimize pandemic burden in terms of case numbers. In future pandemics, where the pathogen behaves similarly to SARS-CoV-2 and no vaccine is available yet but tests are reliable and available in large amounts, we argue that this can be an effective and fast to implement way for policy-makers to keep this impact low. While these measures have been implemented to an extent in the past years in Germany, we believe that a high compliance of the population to these measures is of highest importance to successful mitigation. Thus, sensitizing the population to these measures is key. Model estimations and retrospective results such as the ones presented in this article can be used to accomplish this.
While we have not taken into account mental health and economic impact of these measures, both should be drastically reduced if a full lockdown can be avoided through these measures. Rather than having lasting mobility restrictions for the whole population over a significant time span, only a small portion of persons would have to isolate at a time. While this implication seems natural at first glance, a more detailed behavioral model that includes compliance and mental health would need to be integrated into the whole ABM to analyze NPIs in this regard. The authors of [98] discuss autonomous and policy-induced behavior change during the COVID-19 pandemic, and explain the importance of these aspects in epidemiological models. This field is not explored nearly enough, and will be part of future work. We stress that a reevaluation of NPIs is necessary for a new pathogen, in particular, when its properties vastly differ from the ones considered here, or for different environments, where social infrastructure or cultural habits differ from the setting investigated here. Also, with vaccines being available in a large amount, the importance of testing and isolation can potentially decrease, depending on the political goal. This has not been studied in this article, however.
@@ 5 Conclusion
In this study, we presented an efficiently implemented and parallelized ABM that can be executed with hundreds of thousands to millions of agents on a consumer laptop. Furthermore, it allows the use of a hybrid, shared and distributed memory parallelism to compute large numbers of ensemble results on a supercomputing facility. This capability is essential for handling the computational complexity of ABMs, especially when exploring a wide parameter space or performing sensitivity analysis.
@@
We furthermore studied the combined effects of symptom-based and untargeted, symptom-independent testing strategies, as well as quarantine length and efficiency. We observed that with increased testing of symptomatic persons, the testing of nonsymptomatic persons plays a higher role, leading to a larger reduction in all considered endpoints. On the other hand, with weak symptomatic control, i.e., with a low likelihood to test symptomatic persons, untargeted testing was observed to be mostly inefficient. Furthermore, quarantine length seems to play a more important role than quarantine efficiency and a minimum quarantine of 5–8 days appears to be inevitable for quarantine to have a substantial effect. This, of course, is in line with high viral loads that have been observed at the beginning of the infection. Similar findings have been discussed in the previous section. Finally, we also observed how high symptomatic testing rates leverage quarantining effects, such that even short quarantining can have a large effect on mitigation.
@@ Acknowledgments
The authors gratefully acknowledge computing time on the supercomputer JURECA [73] at Forschungszentrum Jülich under grant HPC4EPI. DK, MMH, and MJK have received funding from the Initiative and Networking Fund of the Helmholtz Association (grant agreement number KA1-Co-08, Project LOKI-Pandemics). The authors DA, CG, AS, and MJK have received funding by the German Federal Ministry for Digital and Transport under grant agreement FKZ19F2211A (Project PANDEMOS), KN and MMH have received funding by the German Federal Ministry for Digital and Transport under grant agreement FKZ19F2211B (Project PANDEMOS). SK has performed his work as part of the Helmholtz School for Data Science in Life, Earth and Energy (HDS-LEE) and received funding from the Helmholtz Association of German Research Centres. The funders had no role in study design, data collection and analysis, decision to publish, or preparation of the manuscript.
@@
Appendix A Comparison of the Munich and Brunswick region
Since data for the urban city of Munich in Germany was readily available, we compared certain trip patterns between Munich and Brunswick to assess whether the findings could be generalized to other urban environments within Germany. Here, we again use the base population of people who live in the respective city plus the incommuters and their households. While the results show a notable similarity between the two cities this comparison is not exhaustive. It provides only an initial insight, as Munich and Brunswick share several key characteristics. However, these findings may not fully apply to smaller cities or non-German urban environments with distinct characteristics, which could exhibit different trip behaviors due to varying socio-economic and cultural factors.
As Fig. 14 illustrates, we have similar age distributions in the synthetic populations of the regions of Brunswick and Munich. In Fig. 15, the heatmap shows the trip frequency between the different location types. We see that both regions share similar trip frequencies between the different location types. Especially, both regions share that most trips go to home location. A similar insight is obtained with Fig. 16. The comparison shows that, relatively, both regions show a very similar distribution of the trip purpose for the overall population.
In Fig. 17, the comparison shows that, relatively, the number of persons for each number of trips is very similar for both regions. Eventually, Fig. 18 also shows that the age-resolved trip data is similar between the two different regions.
Figure 14: Age distribution in the synthetic population of the regions of Brunswick and Munich.
Figure 15: Heatmap of trip frequency of origin-destination type for the locations in the synthetic population of the regions of Brunswick and Munich.
Figure 16: Trip activity distribution for the overall population of in the synthetic population of the regions of Brunswick and Munich.
Figure 17: Number of trips per person in the synthetic population of the regions of Brunswick and Munich.
Figure 18: Trip activity distribution resolved by age groups in the synthetic population of the regions of Brunswick and Munich.